OpenStudy (anonymous):

cot(-105degrees 10'15")

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so you are finding \[\cot(-105^o 10' 15")\] is that correct..?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, so the problem is to convert the angle to units of degrees only, right?

OpenStudy (anonymous):

the directions say determine each value. round to the nearest four decimal places

OpenStudy (anonymous):

Yes, but I'm asking where is your difficulty I assume you may a calculator for cot() but before you can do that you have to convert the degree to angles. do you know how to do that?

OpenStudy (anonymous):

i do not. and I do not have a calculator

OpenStudy (anonymous):

ok, so first thing first. let's convert to degrees. What you see here is degrees + minutes + seconds. Yes I know it's a little confusing because 'minutes' and 'seconds' are time units and not angles. Well, not exactly, they are also angle units. A a single degree divides into 60 'minutes', similar to how hour divides into 60 minutes. A minute divides into 60 seconds, just like with time. So our angle is \(-105^\circ 10'15''\) which means we have a negative angle of 105 degrees, 10 minutes and 15 seconds

OpenStudy (anonymous):

But we know each minute is \(\frac{1}{60}\) degree. so 10 minutes become \(\frac{10}{60}\) degrees which simplifies to what?

OpenStudy (anonymous):

1/6?

OpenStudy (anonymous):

yes. now we have 15 seconds. but a second is \(\frac{1}{60}\) minute, and a minute is \(\frac{1}{60}\) degree.. so a second is really \(\frac{1}{60} \cdot\frac{1}{60} = \frac{1}{3600}\) degree. so 15 seconds will be what?

OpenStudy (anonymous):

15/60

OpenStudy (anonymous):

15 seconds would be 15/60 minutes. but we want to find how many 'degrees' it is. so we say, a minute gets inside a degree 1/60 times. so if we have 15/60 minutes, it's just like 15/60 * 1/60 degrees so that would be 15/(60*60) = 15/3600 Makes sense?

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

ok, simplify that as well, what would that be?

OpenStudy (anonymous):

1/240

OpenStudy (anonymous):

exactly =) So we have: $$ -105^\circ10'15'' = -(105 + \frac{10}{60} + \frac{15}{3600})^\circ = -(105 + \frac{1}{6} + \frac{1}{240})^\circ $$

OpenStudy (anonymous):

Now we have to sum it up and calculate the cot() using a calculator. you said you don't have a calculator but there are plenty of them online: http://web2.0calc.com/

OpenStudy (anonymous):

So first, let's sum it up 105 + 1/6 + 1/240 becomes what?

OpenStudy (anonymous):

0.2711?

OpenStudy (anonymous):

well no. try again

OpenStudy (anonymous):

25241/240?

OpenStudy (anonymous):

That will work too =)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

105.1708

OpenStudy (anonymous):

Ok, good. now we had a negative angle so that is really -105.1708

OpenStudy (anonymous):

now we need to calculate cot() for this angle. make sure the calculator is in mode of degrees and not radians when you do it

OpenStudy (anonymous):

0.2711

OpenStudy (anonymous):

Ye, right =) I thought you gave me that result before for the angle itself, sorry =)

OpenStudy (anonymous):

thanks. can you help me with some more?

OpenStudy (anonymous):

ok, what do you need?

OpenStudy (anonymous):

If f is any function with a period of 6, determine the period of each related function. A: y=f(x+1) B: y=f(1/2 x) C: y=f(3c)+7

OpenStudy (anonymous):

ok, so I assume it is meant to be 3x and not 3c at the last one. let's see Are you familiar with the idea of period?

OpenStudy (anonymous):

yeah sorry it is suppose to be x. i'm kind of familiar with it

OpenStudy (anonymous):

ok, so a periodic function has the exact same behavior every period. that means that no matter what point we pick on the function, if we add or remove one 'period' from that point we get the exact same value. For example, sin() and cos() have period of \(360^\circ\) or \(2\pi\) in radians because no matter what value you pick \(sin(x) = sin(x + 360^\circ)\) By adding 360 you get the exact same value, just a period after.

OpenStudy (anonymous):

makes sense

OpenStudy (anonymous):

So we can see that the exact value of \(x\) doesn't matter. it would work both for x=0 and x=30 in the same way. So in the first example A, we can see that the only difference from the original function f() is that x is incremented by one first. so if it was x=0 now it is x=1. if it was x=5 now it is x=6.. But we already know that it doesn't matter, we just shift all the values of x, but the period is the same.

OpenStudy (anonymous):

We can also examine it. we know that f(x) = f(x+6) because f() has period of 6. Now in A we have $$ y(x) = f(x+1)\\ y(x + 6) = f(x + 6 + 1) = f( (x+1) + 6) = f(x+1) = y(x) $$

OpenStudy (anonymous):

makes sense?